OpenStudy (anonymous):

find equation of the line passing through the points (1, -2) and (-2,4)

OpenStudy (saifoo.khan):

start with the slope.

OpenStudy (anonymous):

ok the slope is my answer?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

then u have to substitue that into a formula to get the answer.

OpenStudy (anonymous):

m=y-y1/x-x2?

OpenStudy (anonymous):

is that the formula?

OpenStudy (saifoo.khan):

\[slope = \frac {y_2-y_1}{x_2-x_1}\]

OpenStudy (precal):

y=-2x is the solution

OpenStudy (saifoo.khan):

after finding the slope, \[({x_2-x_1})m = {y_2-y_1}\] use this.

OpenStudy (anonymous):

ok thanks

OpenStudy (precal):

Better to use y=mx+b to find the y-intercept ie b after the slope formula is used.

OpenStudy (saifoo.khan):

i use the same method.. but most people aren't taught that method in schools. so to make them less confuse i use that same method.